### v2

# The Pound-Drever-Hall Method for Superconducting-Qubit Readout

Abstract
Scaling quantum computers to large sizes requires the implementation of many parallel qubit readouts. Here we present an ultrastable superconducting-qubit readout method using the multi-tone self-phase-referenced Pound-Drever-Hall (PDH) technique, originally developed for use with optical cavities. In this work, we benchmark PDH readout of a single transmon qubit, using room-temperature heterodyne detection of all tones to reconstruct the PDH signal. We demonstrate that PDH qubit readout is insensitive to microwave phase drift, displaying phase stability over 2 hours, and capable of single-shot readout in the presence of phase errors exceeding the phase shift induced by the qubit state. We show that the PDH sideband tones do not cause unwanted measurement-induced state transitions for a transmon qubit, leading to a potential signal enhancement of at least dB over traditional heterodyne readout.
The highest performance optical-frequency measurements often rely on self phase-referenced techniques, such as the Pound-Drever-Hall (PDH) method Drever et al. (1983); Black (2001) for locking lasers and cavities. At lower frequency, and longer wavelength, the path-length stability requirements for phase measurements are less stringent, and many measurement tasks, such as superconducting-qubit readout, are simply carried out by comparing a measurement signal to an external phase reference; however, phase drift remains a serious technical challenge. Here we show that the intrinsically phase-stable PDH technique that originated in the optical frequency domain, and has been used for fast characterization of microwave resonators Lindström et al. (2011); de Graaf et al. (2014); Rath et al. (2025), can be transported to the quantum microwave regime, where it gives rise to improvements for superconducting qubit readout, in terms of long-time stability, scaling complexity, and signal amplification.
Measurement of superconducting qubits Blais et al. (2004, 2021); Wallraff et al. (2005); Koch et al. (2007) is traditionally carried out using dispersive readout, in which the qubit state is inferred from a heterodyne measurement of the state-dependent frequency shifts on a host resonator Blais et al. (2004); Wallraff et al. (2005); Blais et al. (2021). This method has achieved readout speeds below ns and fidelities above 99.7 Swiadek et al. (2024); Spring et al. (2025); Walter et al. (2017). However, heterodyne-based dispersive readout is reliant on high-performance clocking of the microwave generators, AWGs, and digitizers, and generally requires rolling background calibration to compensate for slow drift. The PDH-based superconducting-qubit readout method we present here is not reliant on microwave phase synchronization, and has the potential to improve scalability of superconducting-qubit readout, eliminating the need for rolling reference calibration and potentially reducing the cost per readout channel.
The measured signal of dispersive heterodyne readout after amplification, downconversion, and digital demodulation, takes the form Krantz et al. (2019)
where is the carrier amplitude, is the local oscillator (LO) amplitude, and is the measured phase of the readout signal. In conventional heterodyne readout, the measured phase includes not only the qubit-dependent phase but also contributions from the microwave sources and dispersion in the signal path such that
where is the qubit-dependent phase, and are carrier and LO phase contributions, and is the phase accrued along the propagation path. Fluctuations in depend on the quality of the sources and reference locks, and can be eliminated by modulating a single generator to produce both the measurement and LO tones. However, path-length fluctuations on signals at (e.g., from thermal drifts) remain a source of phase instability even with high-performance references. The resulting phase fluctuations are illustrated in the IQ-plane representation of qubit readout shown in Fig. 1(a). While such fluctuations can be removed through rolling calibration, a natively insensitive method should be more scalable.
Here we explore an approach for detecting qubit-state-dependent frequency shifts on the resonator based on the PDH method. The readout signal in this multi-tone interrogation scheme consists of three copropagating tones: the carrier (), the upper sideband (), and lower sideband (). For phase stability, these three tones are produced from a single source, usually via phase modulation at , and the PDH readout signal, derived from the beat between the carrier and sidebands, takes the form
where and are the amplitudes and phases of the carrier and sidebands. Note: we use the convention that denotes the absolute phase of the lower sideband, which is natural for microwave detection but differs from the standard optical notation by . (See Appendix S2B in the Supplementary Information for details.) In this approach, the signal depends only on differential phases between the carrier and the sidebands, which traverse the same path. Since the sidebands are derived from the same source as the carrier, generator phase errors are common-mode and largely suppressed. Furthermore, unlike in conventional heterodyne readout, where path-length fluctuations comparable to the carrier wavelength produce significant phase shifts, in PDH, only fluctuations comparable to the much longer modulation wavelength matter. Thus, the PDH approach offers an intrinsically phase-stable self-referenced method of detecting the qubit-state-dependent frequency of a microwave resonator, as shown schematically in Fig. 1(b).
In general, large improves readout SNR and speed; however, for weakly anharmonic qubits, such as transmons, large can also induce unwanted excitations Dumas et al. (2024); Connolly et al. (2025); Kurilovich et al. (2025), a process known as measurement-induced state transitions (MIST) Sank et al. (2016); Khezri et al. (2023) or transmon ionization Shillito et al. (2022); Dumas et al. (2024). The PDH technique can potentially mitigate this limitation through intrinsic heterodyne gain, because the PDH signal (Eq. The Pound-Drever-Hall Method for Superconducting-Qubit Readout) scales as . As a result, for sufficiently strong sidebands (), the PDH signal can be much larger for fixed , and can overcome technical noise without adding MIST.
While the PDH technique has been highly successful in the optical regime, full implementation of it in the microwave regime has been limited, primarily due to the lack of high-performance square-law detectors to generate and measure the beatnotes. In this work, we implement a PDH-style multi-tone interrogation of superconducting qubits through simultaneous heterodyne detection of the three tones. Our implementation demonstrates two key features of PDH: (i) the intrinsic phase stability of differential readout, and (ii) the absence of sideband-induced MIST for large sideband offsets.
Phase Stability.— We investigate the long-term phase stability of the two readout approaches by performing PDH-style interrogation. We independently digitize all three tones, comparing the carrier phase, which determines the conventional heterodyne readout signal, with the differential carrier-sideband phase, which determines the PDH signal. These measurements, shown in Fig. 2, illustrate the sensitivity of heterodyne measurement and the robustness of PDH readout. The carrier frequency is fixed near resonance for all datasets, and the modulation frequency varied from MHz to MHz to explore the performance at different sideband offsets. The phases are averaged over 1000 back-to-back single shots and sampled stroboscopically over a period of 2 hours. The carrier phase (Fig. 2(a)) exhibits a linear drift of over that period, due to a frequency mismatch of Hz between the RF and LO generators, i.e. a relative precision error of at a carrier frequency of 6.7 GHz. Although such drift can be suppressed with higher-performance instrumentation and locking schemes (achieving sub Hz offsets and sub-degree drifts below the qubit state phase separation), it illustrates the susceptibility of heterodyne readout to clocking errors. To highlight contributions that are not simply from frequency offsets, the linear component of each trace is subtracted, revealing non-linear fluctuations of up to (Fig. 2(b)), likely caused by temperature drifts or residual phase-locking errors. While improved lab stability or rolling calibration can mitigate these effects, these results emphasize the fundamental sensitivity of heterodyne readout to both clocking errors and laboratory environment.
In contrast, all these sensitivities are removed with PDH readout, in which the sideband is used as a phase reference. The measured differential phases between the carrier and the lower sideband, for sideband detunings between MHz and MHz, are shown in Fig. 2(c). (See Appendix S2 in the Supplementary Information details of the phase convention and extraction procedure.) Despite the absolute drift in , the differential phase remains stable with RMS fluctuations of only , averaged over modulation frequencies. The measured phase stability is independent of sideband detuning even though cryogenic coax typically displays non-negligible dispersion over comparable frequency ranges. This enables the use of large detunings (> MHz), which are expected to be optimal for qubit readout since they allow for large heterodyne gain without added MIST (see Fig. 4) and short detection windows.
In addition to the long-term phase stability measurements above, we demonstrate the single-shot phase stability of PDH and show that a usable signal can be recovered even in the presence of carrier-phase errors comparable to . Measurements are performed with the qubit prepared in both and states and the carrier tuned such that the heterodyne IQ readout separation lies primarily in phase Krantz et al. (2019); Gao et al. (2021). Carrier-phase errors are deliberately induced by unlocking the microwave generators, and the heterodyne measurements corresponding to the two qubit states, shown in Fig. 3(a-i) are essentially indistinguishable. In contrast, PDH readout, shown in Fig. 3(a-ii), preserves separability between the qubit states due to carrier-sideband phase coherence.
A further advantage of PDH readout via simultaneous heterodyne detection of the carrier and sidebands is the possibility of reconstructing more general phase combinations not directly obtainable with a photodiode. In particular, the phase combination , which we term the scissors phase in analogy to the normal modes of ion traps Haroche and Raimond (2006), is orthogonal to the correlated phase shifts induced by errors in both the modulation and detection timing, as well as the generator phase. See Appendix S4 of the Supplementary Information for details. While such timing errors are uncommon in optical PDH setups, where a single radio-frequency source drives both modulation and demodulation, microwave setups often have independent sources. To demonstrate that is robust against modulation and detection errors which compromise PDH, we examine single-shot readout while deliberately inducing timing offsets on all components including the carrier generator, the modulator, and the digitizer. Under these conditions, PDH yields low-fidelity readout (Fig. 3(b-i)) due to quadrature mixing, but the scissors phase, shown in Fig. 3(b-ii), maintains separation of ground and excited states.
Intrinsic Heterodyne Gain.— Because the PDH signal comes from the beats between the carrier and the sidebands, PDH qubit readout can experience effective amplification analogous to mixing with a strong LO if the sidebands are strong compared to the carrier. This intrinsic heterodyne gain can mitigate added noise in down-stream amplifiers and could allow PDH readout with a cryogenic detector to achieve higher readout SNR than conventional heterodyne. Here we demonstrate that sidebands strong enough to achieve more than dB of heterodyne gain can be applied without inducing additional MIST.
To quantitatively benchmark how much sideband amplitude the qubit can tolerate, we evaluate the state transitions versus measurement-tone frequency and power in a two-measurement scheme modified from Ref. Bista et al. (2025), shown schematically in the inset of Fig. 4(a). An initial pulse is applied to the transmon qubit to prepare it in an equal superposition of the ground and excited states. For each configuration, single shots are performed. Each shot consists of two consecutive heterodyne measurements separated by s, and between these, a probe is applied that replicates the effect of a measurement tone with variable-frequency and variable-power. Repeated application of an ideal QND measurement will always yield the same outcome, i.e. the conditional probability . Due to imperfect measurement fidelity, some non-zero transition probabilities ( for ) are always present, but the effect of the probe tone alone can be determined by examining the power and frequency dependence of the transition probabilities, and shown in Fig. 4(a).
To determine the individual measurement outcomes, the IQ plane is divided into three regions, corresponding to the two qubit states, as well as a designated “other” outcome , corresponding to all higher-energy levels of the transmon. Decision boundaries are determined using a reference resonant probe power for which the measurement outcomes are trimodal due to the onset of strong MIST. (See Appendix S6 for details on the decision boundary determination.) Sample IQ histogram plots from the second measurement at MHz detuning are shown in Fig. 4(b), along with the decision boundaries and color-coded outcome regions. All IQ points falling within the same region are assigned to the same state.
The power and detuning dependence of the conditional measurement probabilities are shown in Fig. 4(a). When the probe detuning is zero, the possible range of applied powers is severely limited by the onset of added MIST. Significant increases in and are observed at approximately dBm. See left-hand panel in Fig. 4(a). Above this point, heterodyne readout would be compromised, and the increase of the measurement-induced population with increasing probe power can be clearly seen in the IQ histogram shown in Fig. 4(b-iii).
In the case of PDH readout, however, additional signal amplitude can be gained by leveraging strong off-resonant sidebands. When the detuning of the probe from the carrier frequency exceeds MHz (approximately times the resonator linewidth), no observable signature of added MIST is found within the range of applied probe powers. (See Appendix S6 for additional details and detuning dependence.) Thus, the PDH readout scheme can tolerate at least dBc more power in the sideband at large detunings, corresponding to dB of potential heterodyne gain for the PDH error signal compared to the conventional heterodyne signal, without added MIST.
In conclusion, we have demonstrated that the PDH technique can be successfully extended to the quantum microwave domain as a readout scheme for superconducting qubits, preserving robust phase stability even in the presence of strong dispersion and exhibiting the potential for significant intrinsic heterodyne gain. Our approach, implemented without a cryogenic detector through PDH-style multi-tone interrogation, yields substantial gains in intrinsic phase stability, dramatically reducing both the technical requirements and potential cost per channel of readout electronics, and making the PDH method particularly promising for large-scale integration with superconducting qubits. Ultimately, the development of a Josephson-junction-based cryogenic detector will enable implementation of PDH with intrinsic heterodyne gain, which could surpass conventional heterodyne measurements by achieving higher SNR for fixed resonant probe power. This prospect motivates the development of a new generation of Josephson devices specifically optimized to detect and amplify PDH beatnotes, rather than for linear amplification of the carrier field, paving the way to ultrastable, scalable, and cost-efficient superconducting-qubit readout architectures.
Acknowledgements.
This project was supported by ARL Grants Nos. W911NF-24-2-0107, W911NF-19-2-0181, and W911NF-17-S-0003. IA, WCL, and AK received additional support from NSF Grant No. PHY2047732, the Sloan Fellowship Program, AFOSR Grant No. Grant No. FA9550-21-1-0129, the NSF QLCI grant OMA-2120757, and the University of Maryland. Research at UMD was also supported by the Maryland Quantum Materials Center. We thank Jun Ye, Martin Ritter, and Nathan Schine for helpful comments and discussions, and we thank Jim Phillips and the McRae Group for helpful conversations regarding fast characterization and Q-factor measurements of superconducting resonators using PDH.References
- Drever et al. (1983) R. W. Drever, J. L. Hall, F. V. Kowalski, J. Hough, G. Ford, A. Munley, and H. Ward, Applied Physics B 31, 97 (1983).
- Black (2001) E. D. Black, American journal of physics 69, 79 (2001).
- Lindström et al. (2011) T. Lindström, J. Burnett, M. Oxborrow, and A. Y. Tzalenchuk, Review of Scientific Instruments 82, 104706 (2011).
- de Graaf et al. (2014) S. E. de Graaf, A. Danilov, and S. Kubatkin, IEEE Transactions on Applied Superconductivity 24, 1 (2014).
- Rath et al. (2025) P. K. Rath, J. D. Philips, T. Yoon, K. R. Shirer, A. Fereidouni, and J. Pollanen, arXiv preprint arXiv:2512.06289 (2025).
- Blais et al. (2004) A. Blais, R.-S. Huang, A. Wallraff, S. M. Girvin, and R. J. Schoelkopf, Phys. Rev. A 69, 062320 (2004).
- Blais et al. (2021) A. Blais, A. L. Grimsmo, S. M. Girvin, and A. Wallraff, Rev. Mod. Phys. 93, 025005 (2021).
- Wallraff et al. (2005) A. Wallraff, D. I. Schuster, A. Blais, L. Frunzio, J. Majer, M. H. Devoret, S. M. Girvin, and R. J. Schoelkopf, Physical review letters 95, 060501 (2005).
- Koch et al. (2007) J. Koch, T. M. Yu, J. Gambetta, A. A. Houck, D. I. Schuster, J. Majer, A. Blais, M. H. Devoret, S. M. Girvin, and R. J. Schoelkopf, Phys. Rev. A 76, 042319 (2007).
- Swiadek et al. (2024) F. Swiadek, R. Shillito, P. Magnard, A. Remm, C. Hellings, N. Lacroix, Q. Ficheux, D. C. Zanuz, G. J. Norris, A. Blais, et al., PRX Quantum 5, 040326 (2024).
- Spring et al. (2025) P. A. Spring, L. Milanovic, Y. Sunada, S. Wang, A. F. Van Loo, S. Tamate, and Y. Nakamura, PRX Quantum 6, 020345 (2025).
- Walter et al. (2017) T. Walter, P. Kurpiers, S. Gasparinetti, P. Magnard, A. Potočnik, Y. Salathé, M. Pechal, M. Mondal, M. Oppliger, C. Eichler, and A. Wallraff, Phys. Rev. Appl. 7, 054020 (2017).
- Krantz et al. (2019) P. Krantz, M. Kjaergaard, F. Yan, T. P. Orlando, S. Gustavsson, and W. D. Oliver, Appl. Phys. Rev. 6, 021318 (2019).
- Dumas et al. (2024) M. F. Dumas, B. Groleau-Paré, A. McDonald, M. H. Muñoz Arias, C. Lledó, B. D’Anjou, and A. Blais, Phys. Rev. X 14, 041023 (2024).
- Connolly et al. (2025) T. Connolly, P. D. Kurilovich, V. D. Kurilovich, C. G. Bøttcher, S. Hazra, W. Dai, A. Z. Ding, V. R. Joshi, H. Nho, S. Diamond, et al., arXiv preprint arXiv:2506.05306 (2025).
- Kurilovich et al. (2025) P. D. Kurilovich, T. Connolly, C. G. Bøttcher, D. K. Weiss, S. Hazra, V. R. Joshi, A. Z. Ding, H. Nho, S. Diamond, V. D. Kurilovich, et al., arXiv preprint arXiv:2501.09161 (2025).
- Sank et al. (2016) D. Sank, Z. Chen, M. Khezri, J. Kelly, R. Barends, B. Campbell, Y. Chen, B. Chiaro, A. Dunsworth, A. Fowler, et al., Physical review letters 117, 190503 (2016).
- Khezri et al. (2023) M. Khezri, A. Opremcak, Z. Chen, K. C. Miao, M. McEwen, A. Bengtsson, T. White, O. Naaman, D. Sank, A. N. Korotkov, et al., Physical Review Applied 20, 054008 (2023).
- Shillito et al. (2022) R. Shillito, A. Petrescu, J. Cohen, J. Beall, M. Hauru, M. Ganahl, A. G. Lewis, G. Vidal, and A. Blais, Physical Review Applied 18, 034031 (2022).
- Gao et al. (2021) Y. Y. Gao, M. A. Rol, S. Touzard, and C. Wang, PRX quantum 2, 040202 (2021).
- Haroche and Raimond (2006) S. Haroche and J.-M. Raimond, Exploring the quantum: atoms, cavities, and photons (Oxford university press, 2006).
- Bista et al. (2025) A. Bista, M. Thibodeau, K. Nie, K. Chow, B. K. Clark, and A. Kou, arXiv preprint arXiv:2501.17807 (2025).